### v2

@@ 3.1 Review of the Weyl-$\boldmath Z_{2}$ semimetal
As previously stated, the Weyl-$\boldmath Z_{2}$ semimetal is characterised by the presence of two pairs of nodes with both Weyl and $Z_{2}$ charges. Consequently, the location of these nodes in the momentum space is of significance, as it has the capacity to exert an influence on the topological properties. Depending on this location of nodes, there could be two types of Weyl-$\boldmath Z_{2}$ semimetal, whose topological charges also have slight difference as shown in Fig. 3. In this work, we focus on the left type of Fig. 3 where the two pairs of nodes are located in two directions perpendicular to each other in momentum space and the corresponding topological charges should be (±$(\pm 1,0)$ and $(0,\pm 1)$ for the Weyl and $\boldmath Z_{2}$ charges.
@@ 3.2 Probe fermions on the holographic Weyl-$\boldmath Z_{2}$ semimetal
where $\Gamma_{\pm}:=\frac{1\pm\Gamma^{r}}{2},~{}\hat{\Gamma}_{\pm}:=\frac{1\pm\hat{\Gamma}^{r}}{2}$, $\cdots$ denotes higher order terms, and $k_{1}=(\omega,k_{x},k_{y},k_{z}-q_{1}a_{0}),~{}k_{2}=(\omega,k_{x},k_{y},k_{z}+q_{1}a_{0}),~{}k_{3}=(\omega,k_{x},k_{y}-q_{2}c_{0},k_{z}),~{}k_{4}=(\omega,k_{x},k_{y}+q_{2}c_{0},k_{z})$. It is possible to derive the initial condition of the coupled first-order Dirac equations (3.2) using equation (3.2), as was previously demonstrated in Sec. 2.
@@
The establishment of the Green’s function, $G_{chirality}$ and $G_{\mathbb{Z}_{2}}$, enables the subsequent calculation of the topological Hamiltonian through the implementation of the frequency, $\omega$, set to zero. In the following, the energy spectrum of the topological Hamiltonian will be presented first, and the calculation of the topological invariant, i.e. Weyl charge and $\boldmath Z_{2}$ charge, will be left to Sec. 4.
@@
It is evident that there is a strong similarity between the the chiral and the $\boldmath Z_{2}$ degrees of freedom. We could verify this point in the limit $\frac{M_{1}}{b}\to 0,\frac{M_{2}}{c}\to 0$ by investigating the energy spectrum of the topological Hamiltonian in this limit. The corresponding Green’s functions can be obtained by employing the near horizon solution (3.2), which is
@@
$G_{\boldmath Z_{2}}$ $=-\hat{\Gamma}^{t}\left(\frac{k_{3}^{2m-1}}{4^{m}}\frac{\Gamma\left(\frac{1}{2}-m\right)}{\Gamma\left(\frac{1}{2}+m\right)}k_{3\mu}\hat{\Gamma}^{\mu}\hat{\Gamma}_{+}-\frac{k_{4}^{2m-1}}{4^{m}}\frac{\Gamma\left(\frac{1}{2}-m\right)}{\Gamma\left(\frac{1}{2}+m\right)}k_{4\mu}\hat{\Gamma}^{\mu}\hat{\Gamma}_{-}\right).$ (3.21)
Figure 4: The spectrum function of topological Hamiltonian $H_{Weyl}$(left panel) and $H_{\mathbb{Z}_{2}}$(right panel) at the limit $\frac{M_{1}}{b}\to 0,\frac{M_{2}}{c}\to 0$ with the parameter $b=c=0.8$ for the holographic Weyl $\boldmath{Z}_{2}$ semimetal. Note that the distance between two nodes are determined by the parameter $b$ and $c$.
@@
In the following discussion we will consider the case $\frac{M_{1}}{b},\frac{M_{2}}{c}\neq 0$ in which backreactions have to be taken into consideration, so numerical methods would be employed. Since we have pointed out that when $\frac{M_{1}}{b}=\frac{M_{2}}{c}$ the band structures for $H_{Weyl}$ and $H_{\boldmath Z_{2}}$ will be the same, we will deal with the situation $\frac{M_{1}}{b}\neq\frac{M_{2}}{c}$. Without loss of generality, the value of $\frac{M_{1}}{b}$ is set to be 0.0463 and $\frac{M_{2}}{c}=0.0494$. Then the corresponding spectrums are presented in Fig. 5.
Figure 5: Eight branches of eigenvalues for topological Hamiltonian $H_{Weyl}$(blue line) and $H_{\boldmath Z_{2}}$(red line) at the value $\frac{M_{1}}{b}=0.0463,\frac{M_{2}}{c}=0.0494$. This can be viewed as the band structure of this effective Hamiltonian, therefore we use $E$ to label the vertical axis. Note that all the bands are doubly degenerate. The two nodes denoted by black dots in (b) can be considered as a pair and the the two nodes denoted by black dots in (c) can be considered as a pair which could annihilate into a Dirac point at the critical topological phase transition point.
As demonstrated in Fig. 5, the emergence of multiple nodes is also observed. The distances between a particular pair of nodes exhibit a similar trend when the momentum increases as those found in Fig. 2 for the holographic Weyl semimetal. As there are multiple nodes now, it is important to determine which two nodes should form a pair of nodes with opposite topological charges. A way to examine this is to compare the distances between the nodes as $\frac{M}{b}$ changes and this will be confirmed by the calculation of topological invariant in the next section. When the value of $\frac{M}{b}$ increase up to the critical value $\left(\frac{M}{b}\right)_{c}$, the distance between the two nodes which form a pair will shrink to 00, that is to say, two nodes become a Dirac node at the critical value $\left(\frac{M}{b}\right)_{c}$. Therefore as $\frac{M}{b}$ increases, we can identify with the adjacent two nodes between which the distance decrease as a pair.
@@ 4.1 The spin-Chern number in the condensed matter systems
$H={\bf diag}\left(\begin{pmatrix}\lambda_{SO}+m_{z}&\hbar v_{F}(k_{x}-ik_{y})\\ \hbar v_{F}(k_{x}+ik_{y})&-\lambda_{SO}-m_{z}\end{pmatrix},~{}\begin{pmatrix}-\lambda_{SO}-m_{z}&\hbar v_{F}(k_{x}-ik_{y})\\ \hbar v_{F}(k_{x}+ik_{y})&\lambda_{SO}+m_{z}\end{pmatrix},\right.$
@@
After discussing the example in hexagonal lattice with spin-orbit coupling, it is nature to give the definition for spin Chern number. Given a Hamiltonian $\hat{H}(k)$, if there is a spin operator $\sigma^{z}$ whose eigenvalue is ±$\pm 1$ and it can commute with the Hamiltonian $\hat{H}(k)$, the Hamiltonian can be transformed into a block diagonal matrix under spin operator’s eigenvectors:
@@
$\hat{H}_{\uparrow}$ means that this block can be denoted with the eigenvalue $+1$ for spin operator $\sigma^{z}$, i.e. it corresponds to the spin up part of the system, and $\hat{H}_{\downarrow}$ corresponds to the spin down part respectively. Each block $\hat{H}_{\uparrow},\hat{H}_{\downarrow}$ has their own Weyl charge $C_{\uparrow},C_{\downarrow}$  [^4] In the hexagonal system, one can consider the operator $\mathbb{I}_{2}\otimes\sigma^{3}\otimes\mathbb{I}_{2}$ as the spin operator, so that $C_{\uparrow}=C_{1,1}+C_{-1,1}$ and $C_{\downarrow}=C_{1,-1}+C_{-1,-1}$. In order to shed light on the difference between the spin up and spin down part of the system, the $\boldmath Z_{2}$ topological invariant can be defined as
@@ 4.2.1 The chiral charge
Recall from Fig. 2, we can claim that the nodes for $H_{Weyl}$ only appear on the $k_{z}$ axis and the nodes for $H_{\boldmath Z_{2}}$ only appear on the $k_{y}$ axis. The Chern numbers of several pairs of Weyl nodes have been calculated, and the results demonstrate that each of the Weyl nodes carries the Chern number either $+1$ or $-1$, as is also the case in the weak coupled regime. The explicit values of topological invariants of several nodes calculated from numerics are illustrated in Fig. 6. The nearest nodes denoted with red and green circles have opposite chiral charges and form a pair of nodes which could annihilate in the critical phase transition point to a Dirac semimetal. This has also been confirmed by the behavior of the band structure in the previous section.
Consequently, the numerical outcomes of the topological number demonstrate that when increasing $k_{z}$ from zero, the distances between a particular pair of Weyl nodes initially decrease to (almost) zero and subsequently increase. Without loss of generality, in Fig. 6, the Chern number of the three pairs of nodes around the values of $k_{z}$ equal to 0.48, 1.13 and 1.36, respectively, have been marked (red for $+1$ and green for $-1$). The results demonstrate that, for small $k_{z}$, in each pair of Weyl nodes, the node that has the negative Chern number sits on the left. Conversely, as $k_{z}$ increases, the node that has the positive Chern number is located on the left in place of the one with negative Chern number.
As the system has multiple nodes, the red species of nodes are located in the axis repeatedly and so are the green ones. The distances between the red ones and the distances between the green ones are different so that there is a $k_{zc}\sim 0.76$, at which the red nodes transition from being positioned to the left of the green nodes to being situated to their right. This results in the distance between the paired nodes decreasing with $k_{z}$ at $k<k_{zc}$ and increasing at $k>k_{zc}$. It can thus be posited that the phenomenon of the order of the Chern number changing from $(-1,+1,-1...)$ to $(+1,-1,+1...)$ can be attributed to an interchange between two sets of bands resulting from the competition between two distinct scales: the scale dictating the distance between each red-green node pair and the scale determining the separation among the red nodes themselves. The first scale is determined by $\frac{M}{b}-(\frac{M}{b})_{c}$ and the second scale is determined by the conformal dimension and the charge of the probe fermion. This is a very interesting behavior that is specially found in holographic systems, which we hope to also find in real condensed matter systems.
@@ 4.2.2 The $\boldmath{Z}_{2}$ charge
In Fig. 7, the energy dispersion, determined by $H_{\mathbb{Z}_{2}}$, is plotted as a function of $k_{y}$. The value of the $\boldmath{Z}_{2}$ charge for the three pairs of nodes has also been marked, which also equal to $+1$(read points) or $-1$(green points). In a manner analogous to the phenomenon observed in the chiral degree of freedom, the energy bands undergo an exchange following the attainment of a critical value of $k_{yc}$. The chiral charge is found to be zero for nodes located along the $k_{y}$ axis, while the $\boldmath{Z}_{2}$ also vanish when the nodes located along the $k_{z}$ axis.
In summary, the topological charge for the nodes in the holographic model (3.6) can be expressed as follows: $(1,0),~{}(1,-0),~{}(0,1),~{}(0,-1)$. Within the braket, the first is identified as the chiral charge, and the second as the $\boldmath Z_{2}$ charge. With regard to the configuration depicted on the right side of Fig. 3, the topological charge would be designated as $(1,1),~{}(1,-1),~{}(-1,1),~{}(-1,-1)$. This suggests an alternative topological structure, which will not be further discussed in this paper.
@@ 5 Discussion and Outlook
The topological invariant serves as the fundamental symbol of the distinct topological phases. In this paper, we have calculated the chiral charge and $\boldmath Z_{2}$ charge, which can be treated as the topological invariant of the Weyl/$\boldmath{Z_{2}}$ semimetal in holographic model. By employing four fermions to probe the holographic Weyl/$\boldmath Z_{2}$ semimetal and utilising the topological Hamiltonian approach, the Weyl and $\boldmath Z_{2}$ charges of each Weyl point have been accurately fixed. The Weyl and $\boldmath Z_{2}$ charges, in combination, comprise the corresponding topological invariant. The findings reveal that within the holographic model (3.6), the corresponding topological invariant of the four Weyl nodes are $(1,0),(0,1),(-1,0)$ and $(0,-1)$, respectively.
Furthermore, the existence of multi-Fermi surface structures has been observed in the holographic model, as indicated by calculations of fermion spectrum. The emergence of these multi-Fermi surfaces can be attributed to the phenomenon of strong coupling in holographic model. In turn, the model (3.6) contains multiple pairs of Weyl/$\boldmath Z_{2}$ nodes. We have confirmed which two adjacent Weyl/$\boldmath Z_{2}$ nodes should be a pair of nodes with opposite topological charges that could be annihilated to form a Dirac node at the topological phase transition critical point, both from topological charges and the analysis of the band structures when tuning parameters of the system that determine the distance between each pair of nodes. It is important to highlight that these phenomena are not readily discernible through a mere examination of the $\sigma_{AHE}$ alone. However, further investigation through ARPES may prove beneficial in elucidating this phenomenon.
From another perspective, the Weyl/$\boldmath{Z}_{2}$ semimetal can be conceptualised as a coexistence state between two ideal Weyl semimetals. However, the Weyl/$\boldmath{Z}_{2}$ semimetal exhibits markedly distinct topological properties compared to the ideal Weyl semimetal. Therefore, in order to gain a deeper understanding of the topological phase transition, it is imperative to consider the many-body system which realises the coexistence of disparate topological phases. It has been proposed that a different type of topological many-body system, the topological nodal line semimetal, is capable of coexisting with an ideal Weyl semimetal. Furthermore, a holographic model that realises such a coexisting state has been suggested. In subsequent research, the intention is to calculate the corresponding topological number of the coexisting state, as well as to examine the energy band exchange phenomenon[51].

### v3

@@ 3.1 Review of the Weyl-$\boldmath Z_{2}$ semimetal
As previously stated, the Weyl-$\boldmath Z_{2}$ semimetal is characterised by the presence of two pairs of nodes with both Weyl and $Z_{2}$ charges. Consequently, the location of these nodes in the momentum space is of significance, as it has the capacity to exert an influence on the topological properties. Depending on this location of nodes, there could be two types of Weyl-$\boldmath Z_{2}$ semimetal, whose topological charges also have slight difference as shown in Fig. 3. In this work, we focus on the left type of Fig. 3 where the two pairs of nodes are located in two directions perpendicular to each other in momentum space and the corresponding topological charges should be $(\pm 1,0)$ and $(0,\pm 1)$ for the Weyl and $\boldmath Z_{2}$ charges.
@@ 3.2 Probe fermions on the holographic Weyl-$\boldmath Z_{2}$ semimetal
where $\Gamma_{\pm}:=\frac{1\pm\Gamma^{r}}{2},~\hat{\Gamma}_{\pm}:=\frac{1\pm\hat{\Gamma}^{r}}{2}$, $\cdots$ denotes higher order terms, and $k_{1}=(-\omega,k_{x},k_{y},k_{z}-q_{1}a_{0}),~k_{2}=(-\omega,k_{x},k_{y},k_{z}+q_{1}a_{0}),~k_{3}=(-\omega,k_{x},k_{y}-q_{2}c_{0},k_{z}),~k_{4}=(-\omega,k_{x},k_{y}+q_{2}c_{0},k_{z})$. It is possible to derive the initial condition of the coupled first-order Dirac equations (3.2) using equation (3.2), as was previously demonstrated in Sec. 2.
@@
The establishment of the Green’s function, $G_{chirality}$ and $G_{\boldmath Z_{2}}$, enables the subsequent calculation of the topological Hamiltonian through the implementation of the frequency, $\omega$, set to zero. In the following, the energy spectrum of the topological Hamiltonian will be presented first, and the calculation of the topological invariant, i.e. Weyl charge and $\boldmath Z_{2}$ charge, will be left to Sec. 4.
@@
It is evident that there is a strong similarity between the chiral and the $\boldmath Z_{2}$ degrees of freedom. We could verify this point in the limit $\frac{M_{1}}{b}\to 0,\frac{M_{2}}{c}\to 0$ by investigating the energy spectrum of the topological Hamiltonian in this limit. The corresponding Green’s functions can be obtained by employing the near horizon solution (3.2), which is
@@
[…]
Figure 4: The spectral function of topological Hamiltonian $H_{Weyl}=-{G_{chirality}}^{-1}$, which has been represented in (3.2)(left panel) along $k_{z}$ axis and $H_{\boldmath Z_{2}}=-{G_{\boldmath Z_{2}}}^{-1}$(right panel) along $k_{y}$ axis at the limit $\frac{M_{1}}{b}\to 0,\frac{M_{2}}{c}\to 0$ with the parameter $b=c=0.8$ for the holographic Weyl $\boldmath{Z}_{2}$ semimetal. That is, the eigenvalues for corresponding topological Hamiltonian as a function of $k_{z}$(or $k_{y}$). Note that the distances between two nodes are determined by the parameter $b$ and $c$.
@@
In the following discussion we will consider the case $\frac{M_{1}}{b},\frac{M_{2}}{c}\neq 0$ in which backreaction have to be taken into consideration, so numerical methods would be employed. Since we have pointed out that when $\frac{M_{1}}{b}=\frac{M_{2}}{c}$ the band structures for $H_{Weyl}$ and $H_{\boldmath Z_{2}}$ will be the same, we will deal with the situation $\frac{M_{1}}{b}\neq\frac{M_{2}}{c}$. Without loss of generality, the value of $\frac{M_{1}}{b}$ is set to be 0.0463 and $\frac{M_{2}}{c}=0.0494$. Then the corresponding spectrums are presented in Fig. 5.
Figure 5: Eight branches of eigenvalues for topological Hamiltonian $H_{Weyl}$(blue line) and $H_{\boldmath Z_{2}}$(red line) at the value $\frac{M_{1}}{b}=0.0463,\frac{M_{2}}{c}=0.0494$. This can be viewed as the band structure of this effective Hamiltonian, therefore we use $E$ to label the vertical axis. Note that all the bands are doubly degenerate. The two nodes denoted by black dots in (b) can be considered as a pair and the two nodes denoted by black dots in (c) can be considered as a pair which could annihilate into a Dirac point at the critical topological phase transition point.
As demonstrated in Fig. 5, the emergence of multiple nodes is also observed. The distances between a particular pair of nodes exhibit a similar trend when the momentum increases as those found in Fig. 2 for the holographic Weyl semimetal. As there are multiple nodes now, it is important to determine which two nodes should form a pair of nodes with opposite topological charges. A way to examine this is to compare the distances between the nodes as $\frac{M}{b}$ changes and this will be confirmed by the calculation of topological invariant in the next section. When the value of $\frac{M}{b}$ increases up to the critical value $\left(\frac{M}{b}\right)_{c}$, the distance between the two nodes which form a pair will shrink to 0, that is to say, two nodes become a Dirac node at the critical value $\left(\frac{M}{b}\right)_{c}$. Therefore as $\frac{M}{b}$ increases, we can identify with the adjacent two nodes between which the distance decreases as a pair.
@@ 4.1 The spin-Chern number in the condensed matter systems
[…]
@@
After discussing the example in hexagonal lattice with spin-orbit coupling, it is nature to give the definition for spin Chern number. Given a Hamiltonian $\hat{H}(k)$, if there is a spin operator $\sigma^{z}$ whose eigenvalue is $\pm 1$ and it can commute with the Hamiltonian $\hat{H}(k)$, the Hamiltonian can be transformed into a block diagonal matrix under spin operator’s eigenvectors:
@@
$\hat{H}_{\uparrow}$ means that this block can be denoted with the eigenvalue +1+1 for spin operator $\sigma^{z}$, i.e. it corresponds to the spin up part of the system, and $\hat{H}_{\downarrow}$ corresponds to the spin down part respectively. Each block $\hat{H}_{\uparrow},\hat{H}_{\downarrow}$ has their own Weyl charge $C_{\uparrow},C_{\downarrow}$  [^5] In the hexagonal system, one can consider the operator $\mathbb{I}_{2}\otimes\sigma^{3}\otimes\mathbb{I}_{2}$ as the spin operator, so that $C_{\uparrow}=C_{1,1}+C_{-1,1}$ and $C_{\downarrow}=C_{1,-1}+C_{-1,-1}$. In order to shed light on the difference between the spin up and spin down part of the system, the $\boldmath Z_{2}$ topological invariant can be defined as
@@ 4.2 Topological invariant in the holographic Weyl-$\boldmath{Z}_{2}$ semimetal
Specifically, after obtaining $G_{chirality}$ and $G_{\boldmath Z_{2}}$ numerically for specific $k$, it is possible to get the topological Hamiltonian $H_{Weyl}(k)=-{G_{chirality}}^{-1}(0,k)$, $H_{\boldmath Z_{2}}(k)=-{G_{\boldmath Z_{2}}}^{-1}(0,k)$. To get the Weyl charge and $\boldmath Z_{2}$ charge, we can use the same procedure as in weakly coupled systems, i.e.
  1. 1.
[…]
  2. 2.
    Use the following formula to get the Chern class for each topological Hamiltonian.
[…]
  3. 3.
    Choose the integral surfaces $\Sigma_{Weyl}$ and $\Sigma_{\boldmath Z_{2}}$ which enclose the Weyl node and the $\boldmath Z_{2}$ node respectively, and integrate the Chern class $c_{Weyl}$ and $c_{\boldmath Z_{2}}$ on the integral surfaces to get the corresponding Weyl charge and $\boldmath Z_{2}$ charge
    $Q_{Weyl}=\int_{\Sigma_{Weyl}}c_{Weyl},~Q_{\boldmath Z_{2}}=\int_{\Sigma_{\boldmath Z_{2}}}c_{\boldmath Z_{2}}.$ (4.9)
@@ 4.2.1 The chiral charge
Recall from Fig. 3, we can claim that the nodes for $H_{Weyl}$ only appear on the $k_{z}$ axis and the nodes for $H_{\boldmath Z_{2}}$ only appear on the $k_{y}$ axis. The Chern numbers of several pairs of Weyl nodes have been calculated, and the results demonstrate that each of the Weyl nodes carries the Chern number either +1+1 or 1-1, as is also the case in the weak coupled regime. The explicit values of topological invariants of several nodes calculated from numerics are illustrated in Fig. 6. The nearest nodes denoted with red and green circles have opposite chiral charges and form a pair of nodes which could annihilate in the critical phase transition point to a Dirac semimetal. This has also been confirmed by the behavior of the band structure in the previous section.
Consequently, the numerical outcomes of the topological number demonstrate that when increasing $k_{z}$ from zero, the distances between a particular pair of Weyl nodes initially decrease to (almost) zero and subsequently increase. Without loss of generality, in Fig. 6, the Chern number of the three pairs of nodes around the values of $k_{z}$ equal to 0.48, 1.13 and 1.36, respectively, have been marked (red for +1+1 and green for 1-1). The results demonstrate that, for small $k_{z}$, in each pair of Weyl nodes, the node that has the negative Chern number sits on the left. Conversely, as $k_{z}$ increases, the node that has the positive Chern number is located on the left in place of the one with negative Chern number.
As the system has multiple nodes, the red species of nodes are located in the axis repeatedly and so are the green ones. The distances between the red ones and the distances between the green ones are different so that there is a $k_{zc}\sim 0.76$, at which the red nodes transition from being positioned to the left of the green nodes to being situated to their right. This results in the distance between the paired nodes decreasing with $k_{z}$ at $k<k_{zc}$ and increasing at $k>k_{zc}$. It can thus be posited that the phenomenon of the order of the Chern number changing from (1,+1,1)(-1,+1,-1...) to (+1,1,+1)(+1,-1,+1...) can be attributed to an interchange between two sets of bands resulting from the competition between two distinct scales: the scale dictating the distance between each red-green node pair and the scale determining the separation among the red nodes themselves. The first scale is determined by $\frac{M}{b}-(\frac{M}{b})_{c}$ and the second scale is determined by the conformal dimension and the charge of the probe fermion. This is a very interesting behavior that is specially found in holographic systems, which we hope to also find in real condensed matter systems.
@@ 4.2.2 The $\boldmath{Z}_{2}$ charge
In Fig. 7, the energy dispersion, determined by $H_{\mathbb{Z}_{2}}$, is plotted as a function of $k_{y}$. The value of the $\boldmath{Z}_{2}$ charge for the three pairs of nodes has also been marked, which also equal to +1+1(read points) or 1-1(green points). In a manner analogous to the phenomenon observed in the chiral degree of freedom, the energy bands undergo an exchange following the attainment of a critical value of $k_{yc}$. The chiral charge is found to be zero for nodes located along the $k_{y}$ axis, while the $\boldmath{Z}_{2}$ also vanish when the nodes located along the $k_{z}$ axis.
In summary, the topological charge for the nodes in the holographic model (3.6) can be expressed as follows: (1,0),(1,0),(0,1),(0,1)(1,0),~(1,-0),~(0,1),~(0,-1). Within the braket, the first is identified as the chiral charge, and the second as the $\boldmath Z_{2}$ charge. With regard to the configuration depicted on the right side of Fig. 3, the topological charge would be designated as (1,1),(1,1),(1,1),(1,1)(1,1),~(1,-1),~(-1,1),~(-1,-1). This suggests an alternative topological structure, which will not be further discussed in this paper.
@@ 5 Discussion and Outlook
[…]